OpenStudy (anonymous):

a computer can perform an addition operation in 1.5 x 10^-6 seconds. How many operations can it perform in 1500 seconds?

OpenStudy (aum):

\(1.5 \times 10^{-6}\) seconds for 1 addition. In 1500 seconds how many addition operations? How many \(1.5 \times 10^{-6}\) seconds are there in 1500 seconds? Divide 1500 by \(1.5 \times 10^{-6}\)

OpenStudy (aum):

\[\large \frac{1500}{1.5\times 10^{-6}} = \frac{1.5 \times 10^{3}}{1.5\times 10^{-6}} = \frac{10^3}{10^{-6}} = 10^3 * 10^{6} = 10^{3+6} = 10^9 \]

OpenStudy (aum):

10^9 or a billion operations.